OpenStudy (anonymous):

Given the linear equation y=4/5x-2, find the y-coordinates of the points (-5, )(0, )(5, )... Please show how you got the answer

OpenStudy (anonymous):

y = (4/5)(-5) - 2 = -4 - 2 = -6 y = (4/5)(0) - 2 = 0 - 2 = -2 You should be able to do the last oneee

OpenStudy (anonymous):

first put x = -5 in the given equation y = 4/5(-5) - 2 = -4 -2 = -6 so for first case the missing y coordinate is -6 you can solve the second and third in the same way...